OpenStudy (anonymous):

If a+2b is equal to 125 percent of 4b, what is the value of a/b?

OpenStudy (anonymous):

if you can get a in terms of b (or vice versa) than you can sub it into the 2nd equation. \[a+2b = 1.25(4b)\]\[a=5b-2b\]\[a=3b\] Can you proceed from here?

OpenStudy (anonymous):

Still a bit confused @Mathmuse

OpenStudy (anonymous):

now you sub what 'a' is equal to into the equation:\[\frac{a}{b}=\frac{3b}{b}\] you should be able to simplify this